OpenStudy (anonymous):

Help! Will give you a medal and become a fan (:

OpenStudy (anonymous):

On a separate piece of paper, sketch a unit circle with angle θ in standard position. Use the circle to answer the following questions: For what values of θ is the sine increasing? Decreasing? For what values of θ is the cosine increasing? Decreasing? For which angle between 0° and 360° is sine equal to 0? Where is cosine equal to 0?

OpenStudy (anonymous):

do you have a scientific calculator?

OpenStudy (anonymous):

Yeah, but I have no clue how to work it, it is a TI-84 plus

OpenStudy (anonymous):

so you have to figure out what angle its at when it starts to increase... make your formula first

OpenStudy (anonymous):

Well the angle is going to start at 0 because it is the sine of o. So that must mean that it is going to start increasing as the sine starts going up.

OpenStudy (anonymous):

yes. so your formula would be what?

OpenStudy (anonymous):

well i am confused with the formula, i know it would have to have the sine in it.

OpenStudy (anonymous):

well if it starts at zero, its going to be sine=0x+x

OpenStudy (anonymous):

So everytime you go by up by one or some number you plug it in for x?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So the equation is sin=0x+x or sin = 0x+ y?

OpenStudy (anonymous):

yes so how would you make it when sine is decreasing?

OpenStudy (anonymous):

you would change the + to a -

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

do you think you can finish it?

OpenStudy (anonymous):

Yes, but is it the same thing for cosine?

OpenStudy (anonymous):

i dont think so i dont know cosine very well.

OpenStudy (anonymous):

Ok, i will give it a try! Thanks though for all the help!

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

are you going to have to do more of these kinds of problems?

OpenStudy (anonymous):

Umm yeah, but I am not sure what thequestions are

OpenStudy (anonymous):

Here is the next question Explain how to use the measures of a right triangle to calculate the exact value of sin 30°? How can this information be used to determine the exact value of sin 60°?

OpenStudy (anonymous):

ok well heres some tips : the equation for unit circle is x^2y^2 sin(t)=y cos(t)=x tan(t)=y/x, x is not 0 csc(t)=1/y, y is not 0 sec(t)=1/x, x is not 0 cot(t)=x/y, y is not 0